OpenStudy (anonymous):

Simplify and show work...come so it can be displayed clearly as an equation.

OpenStudy (anonymous):

\[18\sqrt[3]{8}-3\sqrt[3]{27}\]

OpenStudy (anonymous):

the answer is 27, but idk how to get there

OpenStudy (anonymous):

That answer is correct. The cube root of 8 is 2 and the cube root of 27 is 3. Use these values to simplify.

OpenStudy (anonymous):

18(2) - 3(3) = 27

OpenStudy (anonymous):

ohh thanks, haha i totally forgot the cube root of 8 is 2. haha :)